OpenStudy (anonymous):

Simplify: 5x + 2y - 3x + y.

OpenStudy (anonymous):

according to commutative property : a+b +a+c= a+a+b+c hence 5x+2y-3x+y =5x-3x+2y+y =2x+3y

OpenStudy (anonymous):

thanks alot!!!

OpenStudy (anonymous):

my please a lot ,,,,,,,,, medal